# Elements of World Knowledge (EWoK): A cognition-inspired framework for evaluating basic world knowledge in language models

### v1

@@ 4 Evaluation
Using the EWoK framework, we compiled the EWoK-core-1.0 dataset by generating unique fixed substitutions of filler items across templates from domains. In this section, we describe the evaluation of EWoK-core-1.0 via three different evaluation paradigms: traditional plausibility estimates via querying the probability of sentences under the model, LogProbs, as well as two prompt-based strategies, Likert and Choice. The majority of the results reported use the LogProbs evaluation method, which allows comparing a wide range of base and finetuned models. As we show, LogProbs outperforms direct prompting even for large and/or instruction-tuned models.
For the prompt-based evaluations, we collected data from both LLMs as well as human participants using paired identical prompts. Drawing inspiration from comparative psychology, such an approach of matched evaluation has been proposed as a way to support increasingly fair evaluations of LLMs, allowing for more direct comparison of performance, consistency, and context-sensitivity with human participants (Lampinen, 2022).
@@ 4.1 Scoring metrics
For LogProbs evaluation, we use token-level LLM probabilities to calculate as a sum of conditional log probabilities of each token: , where is the vector of tokens composing the target . For Likert, participants (humans and models) are prompted to rate the plausibility of the concatenation of each and pair on a 1–5 scale. For Choice, participants (humans and models) are presented with and , followed by a single target ( or ), and then prompted to select the context ( or ) that better matches the target. For Likert and Choice, details about text generation hyperparameters can be found in A.3 and exact prompt templates can be found in Appendix A.4. The metric for correctness of a given item is the recovery of the designed item structure that and , where score reflects for LogProbs, the integer rating for Likert, and correct context index selection for Choice. In all cases, participants must correctly identify both matches to get the full score ( point). Identification of only one match receives points. In the case of the Likert task, if a model returns the same rating for both pairs, the model receives points. Such a paradigm supports a trivial baseline for all scenarios. Either a random coin flip or a deterministic model generating the same response for each query independent of context will trivially achieve this baseline.
@@ 4.2 Models
We evaluated =20 transformer language models, selected to span a few points in the model design space. Models primarily vary in size (# of parameters; ranging from B–B) and pre-training diet (both # of tokens and source of training corpora). While most evaluated LLMs are dense pre-trained transformers (=13), there are a few one-off comparisons supported including the presence of supervised fine-tuning for instructions (=4) or chat (=2), and mixture-of-experts (MoE) ensembling (=1). We do not intend to draw conclusions about any of these design decisions, but rather to expose variation. Aside from these considerations in exploring variation, the selection of fine-tuned models was filtered to those that did not require specific formatting via the use of a prompt template. Since we evaluate LLMs and humans on identical prompts, it was critical to have complete flexibility in formatting. The full set of evaluated models are listed in Fig. 4 as well as in Appendix A.2.
@@ 4.3 Surface-level item properties
We additionally tested a baseline bag-of-words model based on word2vec embeddings (Mikolov et al., 2013). Embeddings for each word in a context or a target were summed together to derive one vector per context/target. The context/target match was determined by a cosine similarity metric, such that an item is scored correctly if > (and vice versa for ). We then examined the extent to which LLM and human performance correlates with the word2vec baseline. We also tested whether LLM performance correlates with the number of words in an item, as well as with average word frequency in an item using unigram counts from the Google Ngrams (Michel et al., 2011) 2012 American English corpus.
@@ 4.4 Human data
For independent norming of the items in EWoK-core-1.0, we collected data from human participants. We recruited a total of =1,262 participants (591 female, 579 male, 27 other; median age 36; all US-residents with English as their first language) via Prolific, an online study platform. Raters with poor agreement with others () were excluded. The task participants performed was nearly identical to the Likert version of the task for LLMs (in a pilot study, we determined that human results for Likert and Choice show correlation; see Appendix A.1 & B.1 for more details).
5 Release Considerations
Our release-related goals are to (a) reduce the chances of accidental incorporation of EWoK items into LLMs’ training data and (b) promote accountability and reporting when such incorporation is done intentionally. Thus:
EWoK-core-1.0 is released on HuggingFace Datasets (Lhoest et al., 2021) with gated user access to prevent scrapers from accessing it automatically. Users will simply accept a CC-BY license and accompanying Terms of Use (TOU) wherein they will agree to explicitly report any instances when a language model was trained on the EWoK-core-1.0 items, and will be granted access automatically. Link: https://huggingface.co/datasets/ewok-core/EWoK-core-1.0
The code for the EWoK item generation framework is shared in a separate repository on GitHub, with template files downloadable as password-protected archives to prevent automatic scraping. The repository is also protected with a TOU that requires anyone training or fine-tuning on any data generated using EWoK to report that fact. Link: https://github.com/ewok-core/ewok
The code required to replicate the results in this paper, along with human study and model performance data, is shared as a separate GitHub repository following the same protections. Link: https://github.com/ewok-core/ewok-paper.
Although not flawless, this strategy allows us to minimize the chances of accidental incorporation of our data into training sets, as well as to send a clear signal to model developers about the importance of accountable reporting of evaluation dataset use during model training.
@@ 5 Release Considerations
EWoK-core-1.0 is challenging for LLMs
Despite the fact that EWoK-core-1.0 was not designed to be a hard dataset, we found that even the larger models generally perform much below humans: the best model tested, falcon-40b-instruct, has an accuracy of whereas humans perform at (Table AB.2). As expected, larger models tend to do better, although the performance gains are modest. Instruction tuning does not consistently increase or decrease LLM performance under the LogProbs metric.
@@
Figure 4 shows model results by domain, from easiest to hardest (see also Table AB.2). Social interactions is the easiest for both LLMs (mean=, best=) and humans (); spatial interactions is both hardest for LLMs (mean=, best=) and shows the largest gap with human performance (who score highly at ). Overall, domain difficulty is consistent across LLMs, such that hard domains are hard for all and vice versa. One notable exception is the heterogeneous performance of the phi models, with phi-1 consistently among the worst models, phi-1.5 outperforming all models and even humans on physical dynamics, and phi-2 ranging from on par with the largest models on some domains to worse than gpt2-xl on spatial relations. These results can perhaps be attributed to their unique training procedure, which focuses prominently on LLM-generated synthetic data.
@@
Our generation framework is designed to allow easy substitution of different values for names, objects, and locations, among other variables. In principle, these values should not affect the results ( still best matches and not ). However, models show somewhat different performance on the 5 different versions of the dataset, with phi-2 and phi-1.5 showing the largest performance range ( for both). This heterogeneity indicates that arbitrary item choices can substantially affect model performance, such that, in our dataset and in others, “1%” improvements might not be truly meaningful because they would not generalize. Humans show somewhat heterogeneous performance too (range ), although that heterogeneity is driven only by a subset of the domains (Figure 4).
@@
It is critically important to test whether cross-domain performance differences are indeed driven by domain content or are rather attributable to other factors. We consider item design features (direct/indirect context; contrast type for context sentences; and contrast type for target sentences) and surface-level features (average word frequency and sentence length). Figure 5 shows that all these factors affect model performance, often in different ways than they affect humans. Performance of a bag-of-words word2vec-based baseline on different dataset domains and versions is predictive of LLM but not human performance (although baseline performance overall is only ). The number of words in an item negatively affects LLM but not model performance. Unexpectedly, we found a negative relationship between word frequency and both LLM and human performance; follow-up examination of the data showed that this effect is driven by the fact that physical-relations and spatial-relations, the two hardest domains for LLMs, have the highest word frequency. To evaluate relative contributions of these factors to model performance, we jointly modeled all features using mixed effects regression. Design features and surface-level features all contributed to model performance (Table AB.2), with word frequency having a significant positive effect and the number of words having a significant negative effect, as expected. Importantly, domain remained a significant predictor of performance even when accounting for other factors.
@@
Evaluating LLMs with LogProbs resulted in above-chance accuracy for almost all models (Figure 6). As expected, the gap between LogProbs and prompting was larger for smaller models (Hu and Frank, 2024). We used the same prompting setup across all models to assess them in the same way as humans (Lampinen, 2022) (although note that models are also aided via -shot examples). It is possible that targeted prompt-engineering or alternative evaluation strategies will result in above-LogProbs performance, a direction we leave to future work.
Human ratings are often, but not always accurate
Finally, we examined discrepancies between human ratings and experimental labels. Sometimes, the discrepancy resulted from specific fillers changing the plausibility of a pair; for instance The cooler is inside the car. Chao cannot see the cooler. is implausible because the cooler is large and the car has windows, although, for smaller objects and containers without windows, the scenario is plausible. Sometimes, humans made mistakes. One such example is cardinal directions from the spatial relations domain. The scenario The bakery is north of Chao. Chao turns around. The bakery is south of Chao. is implausible because cardinal directions do not depend on the agent’s orientation, and yet our participants often marked it as plausible. Overall, human data collection is a valuable source of information on our dataset but it does not replace ground truth labels.
@@ 7 Discussion
We present a systematic, flexible framework that can be used to test basic world knowledge in language models. Our goal was to develop a dataset that: (1) uses a uniform item format to probe diverse domains of physical and social knowledge, (2) contains items that probe specific concepts (“elements of world knowledge”) within these domains, (3) requires integrating information across sentences, such that the same target sentence is plausible given one context and implausible given another, and (4) consists of generic templates that can be used to generate a wide variety of items.
We then presented evaluation results for a set of openly available models on EWoK-core-1.0, a dataset generated using the EWoK framework. This dataset is moderately challenging for LLMs, with performance varying substantially across domains. We also show that LogProbs contain enough information for most LLMs to perform above chance and that evaluation via prompting underperforms relative to LogProbs even for better-performing models, in agreement with prior results (Hu and Frank, 2024; Hu and Levy, 2023; Hu et al., 2024; Kauf et al., 2024).
@@
The flexibility of our framework allows conducting specific experiments using customized sets of fillers. For instance, one might investigate whether LLMs perform differently on items that include western vs. nonwestern names, items that refer to people by names vs. longer descriptors (“the man in the black hat”), or even items featuring nonwords (like “florp”) instead of real object names.
@@
Knowledge editing research (e.g., Meng et al., 2022, 2023) has often focused on encyclopedic knowledge; but what about knowledge of basic physical and social concepts? Our controlled minimal pair stimuli can allow researchers to identify and manipulate model circuits that might be selectively responsible for knowledge of specific concepts from an array of domains.
@@
For a model to function as a flexible and robust general purpose AI system, it needs to be able to construct, maintain, and update internal world models (Ha and Schmidhuber, 2018; LeCun, 2022) (in cognitive science, variants of such world models are also known as mental models or situation models). The extent to which LLMs possess and use internal world models is subject to ongoing investigation (Hao et al., 2023; Yildirim and Paul, 2024; Wong et al., 2023). The EWoK framework offers an opportunity to combine individual elements of world knowledge to construct multi-step scenarios for evaluating world modeling capabilities in LLMs, within and across physical and social knowledge domains.
@@
Our dataset is written in English; LLM performance might be lower on other languages, especially under-resourced ones. Adapting the EWoK framework to other languages might require redesigning the set of concepts and materials we use, which are currently grounded in the English lexicon. Thus, a multilingual framework can help more cleanly dissociate linguistic and conceptual effects on model performance. Another limitation is that we use the same prompting setup for all models. With tailored prompt engineering, alternative generation methods, or chain-of-thought, LLM performance could improve. Regardless, we expect model generation to conform with general world knowledge under a wide range of evaluation scenarios—the fact that they do not is informative. Finally, due to the synthetic nature of our dataset, some items are semantically weird. Although not necessarily a problem, an alternative is to use LLMs to populate templates. We leave this approach to future work noting that it can introduce confounds that inflate model performance (e.g., Panickssery et al., 2024).

### v2

@@ 4 Evaluation
Using the EWoK framework, we compiled the EWoK-core-1.0 dataset. To do so, our team (which includes experts in specific domains of cognitive science) curated a list of templates that test knowledge of concepts from domains. These templates were then populated with fillers that were randomly sampled from our initial filler set (described in Section 3), resulting in dataset versions. This strategy enables us to explicitly measure the variability associated with the random sampling of filler items (which, in principle, should not affect the results). For EWoK-core-1.0, we chose to hold the set of fillers constant across all items within each version, although that constraint can be relaxed if needed (see Appendix A.1).
We evaluated LLM performance on EWoK-core-1.0 using three different approaches: traditional plausibility estimates via querying the probability of sentences under the model, LogProbs, as well as two prompt-based strategies, Likert and Choice. The majority of the results reported use the LogProbs evaluation method, which allows comparing a wide range of base and finetuned models. As we show, LogProbs outperforms direct prompting even for large and/or instruction-tuned models.
For the prompt-based evaluations, we collected data from both LLMs as well as human participants using paired identical prompts. Drawing inspiration from comparative psychology, such an approach of matched evaluation has been proposed as a way to support increasingly fair evaluations of LLMs, allowing for more direct comparison of performance, consistency, and context-sensitivity with human participants (Lampinen, 2024).
@@ 4.1 Scoring metrics
For LogProbs evaluation, we use token-level LLM probabilities to calculate as a sum of conditional log probabilities of each token: , where is the vector of tokens composing the target . For Likert, participants (humans and models) are prompted to rate the plausibility of the concatenation of each and pair on a 1–5 scale222A 1–5 point scale was chosen to distinguish between degrees of plausibility as opposed to solely absolute direction. Several items in our dataset were distinguishable through this increased granularity, e.g., a 1 vs 2 or 4 vs 5 that would not have been distinguishable otherwise.. This is functionally similar to the LogProbs condition, but explicitly based on the model’s generative behavior as opposed to intrinsic to its scoring. For Choice, participants (humans and models) are presented with and , followed by a single target ( or ), and then prompted to select the context ( or ) that better matches the target. This is another prompting approach that uses the language model discriminatively with access to both items. For Likert and Choice, details about text generation hyperparameters can be found in A.4 and exact prompt templates can be found in Appendix A.5. The metric for correctness of a given item is the recovery of the designed item structure that and , where score reflects for LogProbs, the integer rating for Likert, and correct context index selection for Choice. In all cases, models must correctly identify both matches to get the full score ( point). Identification of only one match receives points. In the case of the Likert task, which has a coarser integer scale, if a model returns the same rating for both pairs, the model receives points. Such a paradigm supports a trivial baseline for all scenarios. Even a random coin flip or a deterministic model generating the same response for each query independent of context will trivially achieve this baseline.
For prompt-based evaluations (Likert and Choice), we used the same prompting setup across all models to assess them in the same way as humans; however, unlike humans, models were additionally provided with -shot examples. These examples did not come from any of our domains and were solely meant to outline the required response formatting including one positive and one negative example. See Appendix A.5 for more details. Our initial experiments revealed that supplying these examples substantially improves model performance, providing us with an opportunity to evaluate their “best shot” at EWoK-core-1.0.
@@ 4.2 Models
We evaluated =20 transformer language models, selected to span a few points in the model design space. Models primarily vary in size (# of parameters; ranging from B–B) and pre-training diet (both # of tokens and source of training corpora). While most evaluated LLMs are dense pre-trained transformers (=13), there are a few one-off comparisons supported including the presence of supervised fine-tuning for instructions (=4) or chat (=2), and mixture-of-experts (MoE) ensembling (=1). We do not intend to draw conclusions about any of these design decisions, but rather to expose variation. Aside from these considerations in exploring variation, the selection of fine-tuned models was filtered to those that did not require specific formatting via the use of a prompt template. Since we evaluate LLMs and humans on identical prompts, it was critical to have complete flexibility in formatting. The full set of evaluated LLMs are listed in Fig. 3 as well as in Appendix A.3.
We additionally tested a baseline bag-of-words model based on word2vec embeddings (Mikolov et al., 2013). Embeddings for each word in a context or a target were summed together to derive one vector per context/target. The context/target match was determined by a cosine similarity metric, such that an item is scored correctly if > (and vice versa for ). Word-embedding-based models provide useful baselines, and there is a long tradition of work in NLI and related tasks utilizing such easy-to-understand baselines to assess task difficulty and the presence of alternative solution strategies Naik et al. (2018). Past NLI work has inadvertently fallen prey to design biases, making it easy for models to rely on heuristics, such as word overlap (McCoy et al., 2020). Additionally, word embeddings can contain rich semantic information (Grand et al., 2022) that may help solve the task by assessing semantic relatedness between words in the context and the target. Using bag-of-words baselines helps understand how robust the task is in the face of such heuristics.
@@ 4.3 Surface-level item properties
To determine the influence of surface-level item properties on model performance, we tested whether LLM performance correlates with the number of words in an item, as well as with average word frequency in an item. Word frequency was determined using unigram counts from the Google Ngrams (Michel et al., 2011) 2012 American English corpus.
@@ 4.4 Human data
For independent norming of the items in EWoK-core-1.0, we collected data from human participants. The task human participants performed was nearly identical to the Likert version of the task for LLMs (in a pilot study, we determined that human results for Likert and Choice show Pearson’s R=0.96 correlation, and decided to drop the Choice condition for the full data collection; see Appendix A.2 (experimental design) & Appendix B.1 (results) for pilot details).
Data collection
To obtain reliable ratings across the full set of EWoK-core-1.0 items, we collected at least 5 responses per item from a total of N=1,262 participants (591 female, 579 male, 27 other/unknown; median age 36; all US-residents who reported English as their first language). Participants were recruited via Prolific (https://prolific.co), an online study platform. To control for quality of data, we excluded 59 participants whose inter-subject Pearson correlation coefficient (compared with average ratings on items from other participant who rated the same items as this subject) was below . Each EWoK-core-1.0 generated item (corresponding to 5 filler-populated versions; see Sec. 3, “Templates & fillers” and Sec. A.1) was split into four subparts: , , , and presented in a Likert setting using the same prompt as that used for LLMs (Appendix A.5; adapted to presentation in a web browser with a free-form text-box for human input). Items were presented so that each participant only saw one of 5 filler-populated variants and one of the four possible sub-items . Therefore, participants provided sensibility judgments independently of any other subparts of the item, closely matching the conditions for LLM evaluation. Each participant rated an average of 57 items (all items came from a single domain). The resultant average inter-subject Pearson correlation across all items was 0.744, where the inter-subject correlation was calculated per-item based on all the participants that rated that item.
Aggregation and evaluation
We average ratings for a given - pair across all (typically, 5) participants that rated the pair. To obtain human norms, we compare the averaged Likert ratings of subparts, awarding point for each half comparison, i.e., and , similar to evaluating LLMs as described in Section 4.1. In case of a tie in scores for a half-comparison, we award 0 points.
@@
EWoK-core-1.0 can be challenging for LLMs
Although EWoK-core-1.0 was not designed to be a difficult dataset, we found that even large open models perform well below human baselines: the best model tested, falcon-40b-instruct, yields a mean accuracy of whereas mean human accuracy is (Table AB.2). As expected, larger models tend to do better, although it is not a sole predictor of performance. Instruction tuning does not consistently increase or decrease LLM performance under the LogProbs metric.
EWoK-core-1.0 was later evaluated on closed frontier models via HELM (Liang et al., 2023) in January of 2025. That evaluation was performed using Choice prompting given that LogProbs for closed models are not available. The highest performing closed models yield performance , which is still below humans () but not by much (see Table 2).
| Model | Mean Accuracy |
| GPT-4 Turbo (2024-04-09) | 0.912 |
| Claude 3.5 Sonnet (2024-06-20) | 0.911 |
| Qwen2 Instruct (72B) | 0.901 |
| Claude 3 Opus (2024-02-29) | 0.893 |
| GPT-4o (2024-05-13) | 0.886 |
| Llama 3 Instruct (70B) | 0.876 |
| GPT-4o mini (2024-07-18) | 0.864 |
| Claude 3 Sonnet (2024-02-29) | 0.848 |
| Mixtral Instruct (8x22B) | 0.842 |
| Mixtral Instruct (8x7B) | 0.837 |
| Claude 3 Haiku (2024-03-07) | 0.829 |
| Claude 3.5 Haiku (2024-10-22) | 0.810 |
| GPT-3.5 Turbo (0125) | 0.797 |
| Mistral Instruct v0.3 (7B) | 0.772 |
| Llama 3 Instruct (8B) | 0.171 |
@@
Figure 3 shows model results by domain, from easiest to hardest (see also Table AB.2). Social interactions is the easiest for both LLMs (mean=, best=) and humans (); spatial interactions is both hardest for LLMs (mean=, best=) and shows the largest performance gap with humans (who score highly at ). We conclude that domains that tap into social knowledge are easier for LLMs whereas domains that require relational knowledge about the physical world are hardest.
Overall, domain difficulty is consistent across LLMs, such that the same domains are predominantly hard or easy for all LLMs. One notable exception is the heterogeneous performance of the phi models, with phi-1 consistently among the worst models, phi-1.5 outperforming all models and even humans on physical dynamics, and phi-2 ranging from on par with the largest models on some domains to worse than gpt2-xl on spatial relations. These results can perhaps be attributed to their unique training procedure, which focuses prominently on LLM-generated synthetic data.
@@
Our generation framework is designed to allow easy substitution of different values for names, objects, and locations, among other variables. In principle, these values should not affect the results ( still best matches and not ). However, models show somewhat different performance on the 5 different versions of the dataset (Table AB.2), with phi-2 and phi-1.5 showing the largest performance range ( for both). This heterogeneity indicates that arbitrary item choices can substantially affect model performance, such that, in our dataset and in others, “1%” improvements might not be truly meaningful because they would not generalize. Humans show somewhat heterogeneous performance too (range ), although that heterogeneity is driven only by a subset of the domains (Figure 3). One of the strengths of EWoK is that it allows us to explicitly explore this variation in details of the dataset construction process. In most benchmarks, these types of details in item-level construction variation are not measurable, even though these effects likely exist.
@@
It is critically important to test whether cross-domain performance differences are indeed driven by domain content or are rather attributable to other factors. We consider item design features (direct/indirect context; contrast type for context sentences; and contrast type for target sentences) and surface-level features (average word frequency and sentence length). Figure 4 shows that all these factors affect model performance, often in different ways than they affect humans. The number of words in an item negatively affects LLM but not human performance. Unexpectedly, we found a negative relationship between word frequency and both LLM and human performance; follow-up examination of the data showed that this effect is driven by the fact that physical-relations and spatial-relations, the two hardest domains for LLMs, have the highest word frequency. To evaluate the relative contributions of these factors to model performance, we jointly modeled all features using mixed effects regression. Design features and surface-level features all contributed to model performance (Table AB.2), with word frequency having a significant positive effect and the number of words having a significant negative effect, as expected. Importantly, domain remained a significant predictor of performance even when accounting for other factors. These results indicate that surface-level factors, such as sentence length and word frequency, contribute but do not fully explain model performance.
Finally, the baseline bag-of-words model, which selects the context with the highest word2vec cosine similarity to the target, performed much worse than LLMs, indicating that simple embedding similarity between words in contexts and targets is insufficient to explain LLM performance. Of note, however, is that the phi-1 model does poorly on all domains and even worse than the baseline bag-of-words model. We believe the overly specific nature of phi-1’s training data that consists of textbook-like demonstrations of coding problems and solutions likely makes it unable to succeed on EWoK-core-1.0. This helps illustrate that without exposure to a naturalistic text corpus which would contain information about the world, even a model that is fluent in following English-language instructions is not able to flexibly reason about the world.
@@
Evaluating LLMs with LogProbs resulted in above-chance accuracy for almost all open models (Figure 5). As expected, the gap between LogProbs and prompting was larger for smaller models (Hu and Frank, 2024). While we observed this general pattern of LogProbs as the dominant strategy, it is possible that highly targeted prompt engineering may result in above-LogProbs performance.
Human ratings are usually, but not always, accurate
Finally, we examined discrepancies between human ratings and experimenter-designed “ground truth” labels. Sometimes, the discrepancy resulted from specific fillers changing the plausibility of a pair; for instance “The cooler is inside the car. Chao cannot see the cooler.” may be implausible because the cooler is a large object and the car has windows, although, for smaller objects and containers without windows, the scenario is more plausible. Sometimes, humans made mistakes. One such example is cardinal directions from the spatial relations domain. The scenario “The bakery is north of Chao. Chao turns around. The bakery is south of Chao.” is implausible because cardinal directions do not depend on the agent’s orientation, and yet our participants often marked it as plausible. Spatial cognition is known to be variable across individuals and cultures (Majid et al., 2004; Pitt et al., 2022), including preferential reliance on relative (‘left’/‘right’) over absolute (‘north’/‘south’) reference frames in Western cultures, which might explain why our (primarily Western) participants performed worse on items with absolute reference frames. Overall, human data collection is a valuable source of information on our dataset but it does not replace ground truth labels.
@@ 6 Discussion
We present a systematic, flexible framework that can be used to test basic world knowledge in language models. Our goal was to develop a dataset that: (1) leverages a uniform item format to probe diverse domains of physical and social knowledge, (2) presents items that employ specific concepts (“elements of world knowledge”) within these domains, (3) requires integrating information across sentences such that the same target sentence is plausible given one context and implausible given another, and (4) consists of generic templates that can be used to generate a large variety of items.
We then presented evaluation results for a set of openly available models on EWoK-core-1.0, a dataset generated using the EWoK framework. This dataset is moderately challenging for LLMs, with performance varying substantially across domains (with social knowledge being the easiest and physical and spatial knowledge being the hardest).
We show that LLM-generated relative item probabilities, estimated with LogProbs allow distinguishing plausible and implausible scenarios above chance for most models and domains and that evaluation via prompting underperforms relative to LogProbs even for better-performing models, in agreement with prior results (Hu and Frank, 2024; Hu and Levy, 2023; Hu et al., 2024; Kauf et al., 2024). Thus, we argue that a natural way to compare models is through the probabilities they output. This is especially the case for models before post-training/instruction-tuning, weak models, or checkpoints, all of which may not have the capability of following complex instructions (e.g., multiple choice) but have substantial knowledge that can be revealed with LogProbs.
We discovered that our dataset has high domain distinguishability—some world knowledge domains are much easier for LLMs than others—but low model distinguishability—many LLMs perform comparably across domains. This pattern indicates that the specifics of the model architecture and training data are less important than the semantic content being evaluated. Results from all models consistently show that social knowledge is easier to learn from text alone than knowledge about the physical world.
We also see that closed frontier models achieve high accuracy on EWoK-core-1.0. This is not surprising, as this dataset was not designed as a challenge benchmark but rather as a comprehensive, broad-coverage dataset of core world knowledge in humans. Thus, the broader framework will have value even when this specific version of the dataset gets saturated.
@@
The flexibility of our framework allows for conducting specific experiments using customized sets of fillers. For instance, one might investigate whether LLMs perform differently on items that include western vs. nonwestern names, items that refer to people by names vs. longer descriptors (“the man in the black hat”), or even items featuring nonwords (like “florp”) instead of real object names.
@@
Knowledge editing research (e.g., Meng et al., 2022, 2023) has often focused on encyclopedic knowledge; but what about knowledge of basic physical and social concepts? Our controlled minimal pair stimuli can allow researchers to identify and manipulate model circuits that might be selectively responsible for knowledge of these specific concepts across several domains.
@@
For a model to function as a flexible and robust general-purpose AI system, it needs to be able to construct, maintain, and update internal world models (Ha and Schmidhuber, 2018; LeCun, 2022) (in cognitive science, variants of such world models are also known as mental models or situation models). The extent to which LLMs possess and use internal world models is subject to ongoing investigation (Hao et al., 2023; Yildirim and Paul, 2024; Wong et al., 2023). The EWoK framework offers an opportunity to combine individual elements of world knowledge to construct multi-step scenarios for evaluating world modeling capabilities in LLMs, within and across physical and social knowledge domains.
@@
Our dataset is written in English; LLM performance might be lower on other languages, especially under-resourced ones. Adapting the EWoK framework to other languages might require redesigning the set of concepts and materials we use, which are currently grounded in the English lexicon. Thus, a multilingual framework can help more cleanly dissociate linguistic and conceptual effects on model performance.
Another limitation is that we use the same prompting setup for all models. With tailored prompt engineering, alternative generation methods, or chain-of-thought reasoning, LLM performance could improve.
Finally, due to the synthetic nature of our dataset, some items may be atypical. We avoid blatant semantic violations by imposing type restrictions on template variables and validating our gold labels against human ratings, so the finalized set of items in EWoK-core-1.0 are semantically valid, just perhaps uncommon. There may be interest in using the EWoK framework to test semantically likely, not just semantically possible items (e.g., Hu et al., 2025). One way to achieve this would be to use LLMs to populate templates. We leave this approach to future work, noting that LLM-based item generation should be used with care to avoid confounds that inflate model performance (e.g., Panickssery et al., 2024a).
Another consequence of template-based generation is that items thus generated may not reflect the distributional statistics of natural language, and thus, log probability scores may not serve as an ideal linking function between the knowledge language models possess and the scenarios we test them on. Methods that explicitly identify ways to tap into LLMs’ reasoning about scenario likelihood may be necessary (e.g., identifying relevant directions in activation-space: Panickssery et al., 2024b). Although we can only explore a limited set of evaluations in a single paper, the EWoK framework is evaluation-method-agnostic, and promises to continue being a useful resource allowing researchers to test for world knowledge in LLMs using novel methods that may be developed in the future.
7 Release Considerations
Our release-related goals are to (a) reduce the chances of accidental incorporation of EWoK items into LLM training data and (b) promote accountability and reporting when such incorporation is done intentionally. Thus:
EWoK-core-1.0 is released on HuggingFace Datasets (Lhoest et al., 2021) with gated user access to prevent scrapers from accessing it automatically. Users will simply accept a CC-BY license and accompanying Terms of Use (ToU) wherein they will agree to explicitly report any instances when a language model was trained on the EWoK-core-1.0 items, and will be granted access automatically. https://huggingface.co/datasets/ewok-core/EWoK-core-1.0
The code for the EWoK item generation framework is shared in a separate repository on GitHub, with template files downloadable as password-protected archives to prevent automatic scraping. The repository is also protected with a ToU that requires anyone training or fine-tuning on any data generated using EWoK to report that fact. https://github.com/ewok-core/ewok
The code required to replicate the results in this paper, along with human study and model performance data, is shared as a separate GitHub repository following the same protections. https://github.com/ewok-core/ewok-paper.